OpenStudy (alexandrad):

You originally draw a design for an art contest on a 2 in. x 5 in. card. The second phase of the contest requires the drawing to be transferred to an 8.5 in x 11 in. standard sheet of paper and utilize as much of the space on the paper as possible. You determine that the largest size one of the dimensions of your drawing can be is 10.5 in. What is the length of the other dimension if the two drawings are similar? Type your exact answer in the blank without the units, and round to the nearest tenths.

OpenStudy (alexandrad):

@zepdrix please help

OpenStudy (alexandrad):

@whpalmer4 please help

OpenStudy (alexandrad):

@1davey29 please help

OpenStudy (alexandrad):

@kidrah69 please help me

OpenStudy (1davey29):

Because 10.5 is larger that 8, it has to be the length of your longer side, which for your note-card is 5. Because the shapes are similar, you can compare their side lengths with a ratio, in this case 2/5=x/10.5. Solve for x, multiply both sides by 10.5, 21/5=x, divide and round, 21/5=4.2, which is already rounded. Your answer, w/o units and rounded, will be 4.2

OpenStudy (alexandrad):

thank you soo much !!!